What is the present worth of a machine with an initial cost of $46,000, an annual cost of $4,000 per year, a salvage value of + $10,000, a useful life of 8 years, and a financing cost rate of 5% per year?
Answer #1

ANSWER:

I = 5% AND N = 8

PW = INITIAL COST + ANNUAL COST(P/A,I,N) + SALVAGE VALUE(P/F,I,N)

PW = - 46,000 - 4,000(P/A,5%,8) + 10,000(P/F,5%,8)

PW = - 46,000 - 4,000 * 6.463 + 10,000 * 0.6768

PW = - 46,000 - 25,852 + 6,768

PW = - 65,084

SO THE PRESENT WORTH IS -$65,084